OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = x4 + 3x3 − 5x2 + 2x − 2, from x = −1 to x = 1.

OpenStudy (anonymous):

-1 1 -5 5

OpenStudy (campbell_st):

well simply plug in x = -1 and evaluate then x = 1and evaluate then the average rate of change is just \[avg..rate..of..change= \frac{f(1) - f(-1)}{1 - (-1)}\] hope it helps

OpenStudy (anonymous):

ok i got -1 for both but that's the wrong answer according to my test. @campbell_st

OpenStudy (campbell_st):

well this is what I got f(1) = -1 f(-1) = -11 perhaps thats helps

OpenStudy (anonymous):

but thats none of the answer choices .@campbell_st

OpenStudy (campbell_st):

well if you use \[\frac{f(1) - f(-1)}{ 1 - (-1)}\] you'll get one of the answers

OpenStudy (anonymous):

it says math processing error. @campbell_st

OpenStudy (campbell_st):

really wow you must be dividing by zero.... which means you have entered it incorrectly... its \[avg..rate.. of..change = (-1 - (-11))\div(1 - (-1))\] enter it the way I've written it

OpenStudy (anonymous):

No no what your entering says "[Math Processing Error]"

OpenStudy (campbell_st):

ok... so its just plan old arithmetic time negative negative is a positive so you have (-1 + 11)/(1 + 1) = that be easy

OpenStudy (anonymous):

OH SO THE ANSWERS 5?

OpenStudy (campbell_st):

thats it.. well done

OpenStudy (anonymous):

Another question , @campbell_st im working on this "Graph the function f(x) = x4 − 7x3 + 12x2 + 4x − 12 using graphing technology and identify for which values of x the graph is increasing."

OpenStudy (campbell_st):

ok... so do you have some technology that graphs the curve..?

OpenStudy (anonymous):

ive graphed it already.

OpenStudy (anonymous):

From x = −3 to x = −2 From x = −2 to x = −1 From x = 1 to x = 2 From x = 2 to x = 3 ^my answers i can choose from

OpenStudy (campbell_st):

ok... so just need to find the parts of the curve where a tangent would have a positive slope... any thought..?

OpenStudy (anonymous):

From x = 1 to x = 2 @campbell_st

OpenStudy (campbell_st):

yep thats what i'd say, well done

OpenStudy (anonymous):

thank you @campbell_st !

OpenStudy (campbell_st):

glad to help